A climbing club planned an expedition. The total cost was $900, which was to be divided equally among the members going. While practicing, three members fell and were injured so they were unable to go. If the cost per person increased by $15, how many people were originally supposed to go?
(If you can show step by step that would be fantastic!)

Answer by jorel555(1290)

Let c be the cost per person for the expedition, and p be the amount of people originally expected to go. Then:
cp=900
p=900/c
and
900/(c+15)=p-3
900/(c+15)=900/c-3
900c=900c+13500-3cē-45c
3cē+45c-13500=0
cē+15c-4500=0
(c+75)(c-60)=0
c=60 or -75
Throwing out the negative number, we get the original cost as $60 per person, which would have been split between 15 people. ☺☺☺☺